# PFGuard: A Generative Framework with Privacy and Fairness Safeguards

### v1

@@ 4.1 Fair Training with Balanced Minibatch Sampling
    We sample $B$ from $D_{i}$ with a resampling probability $w({\mathbf{x}})$ that is proportional to $h({\mathbf{x}})=p_{\text{bal}}({\mathbf{x}})/p_{\text{bias}}({\mathbf{x}})$, which is computed as follows:
    $h({\mathbf{x}})=\frac{p_{\text{bal}}({\mathbf{x}})}{p_{\text{bias}}({\mathbf{x}})}=\frac{p_{\text{bal}}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{\text{bal}}({\mathbf{s}}{=}s)}{p_{\text{bias}}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{\text{bias}}({\mathbf{s}}{=}s)}=\frac{p_{\text{bal}}({\mathbf{s}}{=}s)}{p_{\text{bias}}({\mathbf{s}}{=}s)}\simeq\frac{1/|\mathcal{S}|}{|\{{\mathbf{x}}{\in}D|{\mathbf{s}}{=}s\}|/|D|}$ (5)
    where the second and third equality follows from assumption 1) and 3) above, respectively, and the last approximation follows from assumption 2) above and $D\sim p_{\text{bias}}$.
A sample $B$ from the above procedure is approximately distributed according to $p_{\text{bal}}$ based on SIR; we sample $D_{i}$ from $p_{\text{bias}}$ and resample $B$ from $D_{i}$, where the resampling probability is proportional to $h({\mathbf{x}})=p_{\text{bal}}({\mathbf{x}})/p_{\text{bias}}({\mathbf{x}})$. Since a large number of minibatch samplings is needed to train generative models, the $B$ distribution eventually converges to $p_{\text{bal}}$, leading to a fair generative modeling of $p_{\text{bal}}$.
@@
Our fair sampling technique is also extensible to private settings where the label of sensitive attribute ${\mathbf{s}}$ is unavailable, for example due to privacy regulations (Jagielski et al., 2019; Mozannar et al., 2020; Tran et al., 2022). In such settings, we can employ a binary classification approach to estimate $h({\mathbf{x}})$ like Choi et al. (2020). While their work focuses on non-private settings and assumes an unbiased public reference data on the order of 10%–100% of $|D|$ for the estimation, this assumption can be unrealistic in private domains due to the lack of public data. Our empirical study in Sec. 5.3 shows that we can achieve fairness with only 1–10% of the data, leveraging the ensemble learning of multiple-teacher structure, which can further reduce the estimation error. Note that in this extension, the convergence guarantee may not hold, as $D_{i}\sim p_{\text{bias}}$ in step (2) might not be true in practice if the dataset is randomly partitioned without considering sensitive attribute labels.
@@ 4.2 Private Training with Private Teacher Ensemble Learning
We ensure DP by privatizing knowledge transfer from a teacher ensemble to the generator. Although the sampling technique in Sec. 4.1 transfers more balanced knowledge, privacy risks can be also transferred due to privacy-fairness conflicts. For example, if certain data samples are resampled repeatedly during teacher training, privacy risks like memorization can occur in the teacher models and be transferred to the generator. We thus privatize the knowledge transfer with DP techniques to provide strict DP guarantees in the generator. Note that only the generator needs to have privacy as it is the one that is released publicly to produce synthetic data.
@@
We utilize Private Teacher Ensemble Learning (PTEL) (Papernot et al., 2016; 2018) to ensure DP in the knowledge transfer. Compared to non-private ensemble learning, PTEL 1) assumes each teacher model is trained on a disjoint data subset and 2) adds noise proportional to the sensitivity of the knowledge aggregation. Here, sensitivity is derived from data disjointness, where a single data point affects at most one teacher. For example, GNMax aggregator (Papernot et al., 2018) aggregates prediction of teacher models $\{T_{i}\}^{n_{T}}_{i=1}$ on a data sample ${\mathbf{x}}$ for its class label as follows:
$\text{GNMax}({\mathbf{x}})=\operatorname*{arg\,max}_{j}\{n_{j}({\mathbf{x}})+\mathcal{N}(0,\sigma^{2})\}\quad\text{for}~{}j=1,...,c$ (6)
where $n_{j}({\mathbf{x}})$ denotes the vote count for the $j$-th class (i.e., $n_{j}({\mathbf{x}})=|\{i:T_{i}({\mathbf{x}}){=}j\}|$), and $\mathcal{N}(0,\sigma^{2})$ denotes random Gaussian noise. Here, the $l_{2}$-sensitivity (Def. 2.2) is $\sqrt{2}$, as a single data point affects at most one teacher, increasing the vote counts by 1 for one class and decreasing the count by 1 for another class (see a more detailed analysis in Sec. B.2). Consequently, the GNMax aggregator satisfies $(\varepsilon,\delta)$-DP for $\sigma\geq\sqrt{8\ln(1.25/\delta)}/\varepsilon$ based on the Gaussian mechanism (Thm. 2.1).
@@
PFGuard can be easily integrated with existing PTEL-based generative models by simply modifying the minibatch sampling process as described in Sec. 4.1. PTEL has been widely adopted to train generators with privatized teacher output to ensure DP (Jordon et al., 2018; Chen et al., 2020; Long et al., 2021; Wang et al., 2021a). Although the exact sensitivity values of these PTEL-based generative models vary depending on what teacher knowledge is aggregated (e.g., votes on class labels (Jordon et al., 2018) or gradient directions (Wang et al., 2021a)), PFGuard preserves any sensitivity as long as the PTEL enforce data disjointness; even with fair sampling, a single data point still affects only one teacher. PFGuard is thus compatible with various PTEL-based generative models, enhancing fairness while preserving DP guarantees – see Sec. C.1 for the full algorithm.
@@ 4.3 Advantages of Integrated Design
PFGuard can sidestep privacy breaches and fairness cancellation arising from privacy-fairness conflicts. Applying fairness-only techniques to existing DP generators can compromise DP guarantees and require complex sensitivity assessments; in contrast, PFGuard automatically preserves DP guarantees of any PTEL-based DP generators through data disjointness, eliminating the need of such assessments. Privacy-only techniques often use sample-level bounding, which directly limits an individual sample’s influence (e.g., gradient clipping discussed in Sec. 3) and can lead to fairness cancellation by suppressing the use of minority data. In contrast, PFGuard uses indirect privacy bounding, in the sense that we limit the knowledge transfer of teacher models in order to limit individual sample’s influence. Since there are no DP constraints during the teacher learning, the teacher models can effectively learn balanced knowledge across data groups.
@@
The fair training of PFGuard adds minimal training complexity, preserving the utility for the subsequent private training stage. The proposed sampling technique requires a simple modification in the minibatch sampling process for fairness, avoiding the need for additional fairness loss terms (Sattigeri et al., 2019; Yu et al., 2020) or auxiliary classifiers (Tan et al., 2020; Um & Suh, 2023) typically employed in fairness-only techniques. In Sec. 5, we also show that PFGuard incurs negligible overhead in computation time when integrated with existing PTEL-based generative models.
@@ 5 Experiments
Datasets     We evaluate PFGuard on three image datasets: 1) MNIST (LeCun et al., 1998) and FashionMNIST (Xiao et al., 2017) for various analyses and baseline comparisons, and 2) CelebA (Liu et al., 2015) to observe performance in real-world scenarios more closely related to privacy and fairness concerns. Here, MNIST contains handwritten digit images, FashionMNIST contains clothing item images, and CelebA contains facial images. While MNIST and FashionMNIST are simplistic and less reflective of real-world biases, they enable reliable fairness analyses on top of high-performing DP generative models on these datasets, making them widely adopted in recent studies addressing the privacy-fairness intersections (Bagdasaryan et al., 2019; Farrand et al., 2020; Ganev et al., 2022). For CelebA, we resize the images to 32 $\times$ 32 $\times$ 3 (i.e., CelebA(S)) and to 64 $\times$ 64 $\times$ 3 (i.e., CelebA(L)) following the conventions in the DP generative model literature (Long et al., 2021; Wang et al., 2021a; Cao et al., 2021). More dataset details are in Sec. D.1.
Bias Settings     We create various bias settings across classes and subgroups, focusing on four scenarios: 1) binary class bias, which is a basic scenario often addressed in DP generative models, and 2) multi-class bias, subgroup bias, and unknown subgroup bias, which are more challenging scenarios typically addressed in fairness techniques, but not in DP generative models. We observe that DP generative models mostly perform poorly in these challenging scenarios, especially with complex datasets like CelebA, so we use MNIST for more reliable analyses. While recent privacy-fairness studies on MNIST mostly focus on class bias (Bagdasaryan et al., 2019; Farrand et al., 2020), we additionally analyze subgroup bias using image rotation for more fine-grained fairness analyses and to support prominent fairness metrics like equalized odds (Hardt et al., 2016). In all experiments, we denote $Y=0$ as the minority class and $S=0$ as the minority group. More details on bias levels (e.g., size ratios between majority and minority data) and bias creation are in Sec. D.1.
@@
  • Fairness. We measure the group size disparity in synthetic data with the KL divergence to uniform distribution $U(S)$ (i.e., $D_{KL}(p_{G}(S)||U(S))$(Yu et al., 2020) and the distribution disparity (i.e., $|p_{G}(S)-U(S)|$(Choi et al., 2020), where $p_{G}(S)$ denotes generated distribution w.r.t. $S$. We measure the fairness disparities in downstream tasks as follows: equalized odds disparity (i.e., $\max_{y,s_{1},s_{2}}|\Pr(\hat{Y}{=}y|Y{=}y,S{=}s_{1}){-}\Pr(\hat{Y}{=}y|Y{=}y,S{=}s_{2})|,~{}\forall y\in\mathcal{Y}$, $s_{1},s_{2}{\in}\mathcal{S}$), demographic disparity (i.e., $\max_{s_{1},s_{2}}|\Pr(\hat{Y}{=}1|S{=}s_{1}){-}\Pr(\hat{Y}{=}1|S{=}s_{2})|,~{}\forall s_{1},s_{2}\in\mathcal{S}$), and accuracy disparity (i.e., $\max_{y_{1},y_{2}}|\Pr(\hat{Y}{=}y_{1}|Y{=}y_{1}){-}\Pr(\hat{Y}{=}y_{2}|Y{=}y_{2})|,~{}\forall y_{1},y_{2}\in\mathcal{Y}$).
  • Privacy. We use privacy budget $\varepsilon$ for DP (Def. 2.1), which is preserved in both synthetic data and downstream tasks due to the post-processing property of DP (see more details in Sec. B.1).
@@ 5.2 Privacy-Fairness-Utility Tradeoff
We compare our privacy-fairness-utility performance with naïve combinations of prior approaches. We evaluate performance under two bias settings: 1) subgroup bias and 2) unknown subgroup bias. Table 5.2 shows the results, which aligns with our privacy-fairness counteraction analysis in Sec. 3. On the one hand, fairness-only reweighting approaches compromise privacy due to the increased iterations from modifying the loss function for fair training (i.e., the more a model uses the data, the weaker privacy it provides). On the other hand, privacy-fairness classification techniques maintain the original privacy guarantees, but significantly degrade utility and fairness, resulting in lower image quality and size disparities across groups. We further discuss this fairness-utility degradation in Sec. A, where we suspect that directly changing gradient clipping thresholds for fairness may severely affect utility when used in generative settings. In contrast, PFGuard is the only method that successfully achieves both privacy and fairness and preserves the closest utility to the original models.
@@ 5.3 Ablation Study
We validate the proposed theoretical upper bound on the number of teachers for fairness, which depends on the bias level of the training data. To effectively simulate scenarios where a teacher receives only a small subset of minority data, we evaluate PFGuard in a multi-class bias setting, downsizing the minority class (i.e., class 8 for MNIST) by a factor of $\gamma$. Given that MNIST has fewer than 6,000 samples for class 8, our proposed upper bound is $\gamma{\leq}5$ if we fix the number of teachers to 1,000. Fig. 5 shows that exceeding $\gamma{=}5$ leads to a noticeable decline in accuracy for the minority class, which is consistent with our theoretical results. It is noteworthy that even with the decline, PFGuard shows higher accuracy than the privacy-only baseline, which shows a consistent decrease in accuracy for the minority as $\gamma$ increases.
@@
We explore the influence of the reference dataset size when PFGuard is extended to unknown subgroup bias setting. Fig. 5 shows PFGuard achieves comparable fairness even with a small reference dataset size, while showing a slight decline in the overall utility.
More Analyses We provide more experiments in Sec. E, including a comparison of computation time (Sec. E.1), results on different datasets such as FashionMNIST (Sec. E.2), and employing an additional normalization technique to further enhance the overall image quality (Sec. E.3).

### v2

@@ 4.1 Fair Training with Balanced Minibatch Sampling
    We sample $\mathcal{B}$ from $\mathcal{D}_{i}$ with a resampling probability $w({\mathbf{x}})$ that is proportional to $h({\mathbf{x}})=p_{\text{bal}}({\mathbf{x}})/p_{\text{bias}}({\mathbf{x}})$, which is computed as follows:
    $h({\mathbf{x}})=\frac{p_{\text{bal}}({\mathbf{x}})}{p_{\text{bias}}({\mathbf{x}})}=\frac{p_{\text{bal}}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{\text{bal}}({\mathbf{s}}{=}s)}{p_{\text{bias}}({\mathbf{x}}|{\mathbf{s}}{=}s)p_{\text{bias}}({\mathbf{s}}{=}s)}=\frac{p_{\text{bal}}({\mathbf{s}}{=}s)}{p_{\text{bias}}({\mathbf{s}}{=}s)}\simeq\frac{1/|\mathcal{S}|}{|\{{\mathbf{x}}{\in}\mathcal{D}|{\mathbf{s}}{=}s\}|/|\mathcal{D}|}$ (5)
    where the second and third equality follows from assumption 1) and 3) above, respectively, and the last approximation follows from assumption 2) above and $\mathcal{D}\sim p_{\text{bias}}$.
A sample $\mathcal{B}$ from the above procedure is approximately distributed according to $p_{\text{bal}}$ based on SIR; we sample $\mathcal{D}_{i}$ from $p_{\text{bias}}$ and resample $\mathcal{B}$ from $\mathcal{D}_{i}$, where the resampling probability is proportional to $h({\mathbf{x}})=p_{\text{bal}}({\mathbf{x}})/p_{\text{bias}}({\mathbf{x}})$. Since a large number of minibatch samplings is needed to train generative models, the $\mathcal{B}$ distribution eventually converges to $p_{\text{bal}}$, leading to a fair generative modeling of $p_{\text{bal}}$.
@@
Our fair sampling technique is also extensible to private settings where the label of sensitive attribute ${\mathbf{s}}$ is unavailable, for example due to privacy regulations (Jagielski et al., 2019; Mozannar et al., 2020; Tran et al., 2022). In such settings, we can employ a binary classification approach to estimate $h({\mathbf{x}})$ like Choi et al. (2020). While their work focuses on non-private settings and assumes an unbiased public reference data on the order of 10%–100% of $|\mathcal{D}|$ for the estimation, this assumption can be unrealistic in private domains due to the lack of public data. Our empirical study in Sec. 5.3 shows that we can achieve fairness with only 1–10% of the data, leveraging the ensemble learning of multiple-teacher structure, which can further reduce the estimation error. Note that in this extension, the convergence guarantee may not hold, as $\mathcal{D}_{i}\sim p_{\text{bias}}$ in step (2) might not be true in practice if the dataset is randomly partitioned without considering sensitive attribute labels.
@@ 4.2 Private Training with Private Teacher Ensemble Learning
Although the fair sampling technique in Sec. 4.1 can enhance fairness, it can also introduce privacy risks due to inherent privacy-fairness conflicts (Sec. 3). Specifically, repeated resampling of certain data samples during teacher training can lead to memorization, compromising privacy in teacher models and ultimately affecting the generator during knowledge transfer. To address this privacy risk, we privatize knowledge transfer rather than the teacher models themselves, ensuring DP solely in the generator. Since the generator is only needed after training for synthetic data generation, we can only release DP generator to the public (Chen et al., 2020; Long et al., 2021).
@@
We utilize Private Teacher Ensemble Learning (PTEL) (Papernot et al., 2016; 2018) to ensure DP during knowledge transfer. Unlike non-private ensemble learning, PTEL 1) trains each teacher model on a disjoint data subset and 2) adds noise proportional to the sensitivity of the aggregated knowledge (e.g., class labels (Papernot et al., 2018), gradients (Chen et al., 2020), etc.). Although aggregated knowledge can differ, sensitivity is commonly derived from data disjointness, where a single data point affects at most one teacher. For instance, GNMax aggregator (Papernot et al., 2018) aggregates predicted class label for a query input $\bar{{\mathbf{x}}}$ from teachers $\{T_{i}\}^{n_{T}}_{i=1}$ as follows:
$\text{GNMax}(\bar{{\mathbf{x}}})=\operatorname*{arg\,max}_{j}\{n_{j}(\bar{{\mathbf{x}}})+\mathcal{N}(0,\sigma^{2})\}\quad\text{for}~{}j=1,...,c$ (6)
where $n_{j}$ denotes the vote count for the $j$-th class (i.e., $n_{j}(\bar{{\mathbf{x}}})=|\{i:T_{i}(\bar{{\mathbf{x}}}){=}j\}|$), and $\mathcal{N}(0,\sigma^{2})$ denotes random Gaussian noise. Here, the $l_{2}$-sensitivity (Def. 2.2) is $\sqrt{2}$, as a single data point affects at most one teacher, increasing the vote counts by 1 for one class and decreasing the count by 1 for another class (see a more detailed analysis in Sec. B.2). Consequently, the GNMax aggregator satisfies $(\varepsilon,\delta)$-DP for $\sigma\geq\sqrt{8\ln(1.25/\delta)}/\varepsilon$ based on the Gaussian mechanism (Thm. 2.1).
@@
PFGuard edits existing PTEL-based generative models to achieve both privacy and fairness by simply modifying the minibatch sampling process as described in Sec. 4.1. PTEL has been widely used to train DP generators (Jordon et al., 2018; Chen et al., 2020; Long et al., 2021; Wang et al., 2021a), resulting in various sensitivity analyses. While integrating fair training techniques can invalidate these analyses as discussed in Sec 3, PFGuard preserves any sensitivity value as long as PTEL enforce data disjointness; even with fair sampling, a single data point still affects only one teacher. PFGuard thus enhances fairness of various PTEL-based generative models, while preserving their own DP guarantees. We present formal DP guarantees and the training algorithm in Sec. C.
@@ 4.3 Advantages of Integrated Design
PFGuard can sidestep privacy breaches and fairness cancellation arising from privacy-fairness conflicts. Compared to existing fairness-only techniques, which can compromise DP guarantees and require complex sensitivity assessments when applied to DP generators, PFGuard automatically preserves DP guarantees of PTEL-based DP generators by maintaining sensitivities, eliminating the need of complex assessments. Compared to existing privacy-only techniques, which often directly limit an individual sample’s influence (e.g., gradient clipping discussed in Sec. 3) and can lead to fairness cancellation by suppressing the use of minority data, PFGuard uses indirect privacy bounding, in the sense that we limit the influence of teacher models in order to limit individual sample’s influence. Since there are no DP constraints during the teacher learning, the teacher models can effectively learn balanced knowledge across data groups.
@@
The fair sampling of PFGuard introduces minimal training complexity, which can preserve utility while enhancing fairness. By modifying only the minibatch sampling process, PFGuard maintains the original loss function and avoids the additional fairness loss terms (Sattigeri et al., 2019; Yu et al., 2020) or auxiliary classifiers (Tan et al., 2020; Um & Suh, 2023) typically employed in fairness techniques. As a result, we show how PFGuard can achieve a prominent fairness-utility tradeoff with negligible computational overhead when integrated into PTEL-based generative models in Sec. 5. Further discussions on PFGuard’s 1) flexibility with other fairness techniques and 2) synergy with adversarial learning are presented in Sec. C.3.
@@ 5 Experiments
Datasets     We evaluate PFGuard on three image datasets: 1) MNIST (LeCun et al., 1998) and FashionMNIST (Xiao et al., 2017) for various analyses and baseline comparisons, and 2) CelebA (Liu et al., 2015) to observe performance in real-world scenarios more closely related to privacy and fairness concerns. MNIST contains handwritten digit images, FashionMNIST contains clothing item images, and CelebA contains facial images. While MNIST and FashionMNIST are simplistic and less reflective of real-world biases, they enable reliable fairness analyses on top of high-performing DP generative models on these datasets, making them widely adopted in recent studies addressing the privacy-fairness intersections (Bagdasaryan et al., 2019; Farrand et al., 2020; Ganev et al., 2022). For CelebA, we resize the images to 32 $\times$ 32 $\times$ 3 (i.e., CelebA(S)) and to 64 $\times$ 64 $\times$ 3 (i.e., CelebA(L)) following the conventions in the DP generative model literature (Long et al., 2021; Wang et al., 2021a; Cao et al., 2021). We provide more dataset details (e.g., dataset sizes) in Sec. D.1.
Bias Settings     We create various bias settings across classes and subgroups, focusing on four scenarios: 1) binary class bias, which is a basic scenario often addressed in DP generative models, and 2) multi-class bias, subgroup bias, and unknown subgroup bias, which are more challenging scenarios typically addressed in fairness techniques, but not in DP generative models. We observe that DP generative models mostly perform poorly in these challenging scenarios, especially with complex datasets like CelebA, so we use MNIST for more reliable analyses. While recent privacy-fairness studies primarily focus on class bias in MNIST (Bagdasaryan et al., 2019; Farrand et al., 2020), we additionally analyze subgroup bias using image rotation for more fine-grained fairness analyses and to support prominent fairness metrics like equalized odds (Hardt et al., 2016). In all experiments, we denote ${\mathbf{y}}=0$ as the minority class and ${\mathbf{s}}=0$ as the minority group. More details on bias levels (e.g., size ratios between majority and minority data) and bias creation are in Sec. D.1.
@@
  • Fairness. We measure the group size disparity in synthetic data with the KL divergence to uniform distribution $U({\mathbf{s}})$ (i.e., $D_{KL}(p_{G}({\mathbf{s}})||U({\mathbf{s}}))$(Yu et al., 2020) and the distribution disparity (i.e., $|p_{G}({\mathbf{s}})-U({\mathbf{s}})|$(Choi et al., 2020), where $p_{G}({\mathbf{s}})$ denotes generated distribution w.r.t. ${\mathbf{s}}$. We measure the fairness disparities in downstream tasks as follows: equalized odds disparity (i.e., $\max_{y,s_{1},s_{2}}|\Pr(\hat{{\mathbf{y}}}{=}y|Y{=}y,{\mathbf{s}}{=}s_{1})-\Pr(\hat{{\mathbf{y}}}{=}y|{\mathbf{y}}{=}y,{\mathbf{s}}{=}s_{2})|,~{}\forall y\in\mathcal{Y}$, $s_{1},s_{2}{\in}\mathcal{S}$), demographic disparity (i.e., $\max_{s_{1},s_{2}}|\Pr(\hat{{\mathbf{y}}}{=}1|{\mathbf{s}}{=}s_{1})-\Pr(\hat{{\mathbf{y}}}{=}1|{\mathbf{s}}{=}s_{2})|,~{}\forall s_{1},s_{2}\in\mathcal{S}$), and accuracy disparity (i.e., $\max_{y_{1},y_{2}}|\Pr(\hat{{\mathbf{y}}}{=}y_{1}|{\mathbf{y}}{=}y_{1})-\Pr(\hat{{\mathbf{y}}}{=}y_{2}|{\mathbf{y}}{=}y_{2})|,~{}\forall y_{1},y_{2}\in\mathcal{Y}$).
  • Privacy. We use privacy budget $\varepsilon$ for DP (Def. 2.1), which is preserved in both synthetic data and downstream tasks due to the post-processing property of DP (see Sec. B.1 for more details).
@@ 5.2 Privacy-Fairness-Utility Tradeoff
We compare our privacy-fairness-utility performance against simple combinations of prior approaches. We evaluate performance under two bias settings: 1) subgroup bias and 2) unknown subgroup bias. Table 5.2 shows the results, which aligns with our privacy-fairness counteraction analysis in Sec. 3; fairness-only reweighting approaches compromise privacy due to the increased iterations, which may arise from modifying the loss function for fair training (i.e., the more a model uses the data, the weaker privacy it provides). In comparison, privacy-fairness classification techniques can maintain the original privacy guarantees, but significantly degrade utility and fairness, resulting in lower image quality and size disparities across groups – further discussions are provided in Sec. A. Among all methods, PFGuard is the only method that successfully 1) achieves both privacy and fairness and 2) preserves the closest utility to the original models. In Sec. E.1, we provide more analyses of the privacy-fairness-utility tradeoff, including Pareto Frontier results and varying privacy levels.
@@ 5.3 Ablation Study
We validate the proposed theoretical upper bound on the number of teachers for fairness, which depends on the bias level of the training data. To effectively simulate scenarios where a teacher receives only a small subset of minority data, we evaluate PFGuard in a multi-class bias setting, downsizing the minority class (i.e., class ‘8’ for MNIST) by a factor of $\gamma$. Given that MNIST has fewer than 6,000 samples for class ‘8’, our proposed upper bound is $\gamma{\leq}5$ if we fix the number of teachers to 1,000. Fig. 5 shows that exceeding $\gamma{=}5$ leads to a noticeable accuracy drop for the minority class, which is consistent with our theoretical results. It is noteworthy that even with the decline, PFGuard shows higher accuracy than the privacy-only baseline, which shows a consistent decrease in accuracy for the minority as $\gamma$ increases.
@@
We explore the influence of the reference dataset size when PFGuard is extended to unknown subgroup bias setting. Fig. 5 shows PFGuard achieves comparable fairness even with a small reference dataset size, while showing a slight increase in the overall FID.
More Analyses We provide more experiments in Sec. E, including a comparison of computation time (Sec. E.3), results on different datasets (Sec. E.2 and E.7), performance with additional normalization techniques (Sec. E.4), and a performance comparison with a different sampling strategy (Sec. E.5).